OpenStudy (anonymous):

how do you do 3(x-1)^2/3 +4 =52 you have to check for extraneous solutions

OpenStudy (anonymous):

first leave only (x-1)^2/3 at left hand side (LHS). next u square root both sides. then u multiply both sides to power of 3. after that u rearrange it to get x.

OpenStudy (anonymous):

\[3(x-1)^{2/3}=48 \]

OpenStudy (anonymous):

\[(x-1)^{2/3}=16\]

OpenStudy (anonymous):

\[((x-1)^{2/3})^{3/2}=16^{3/2}\]

OpenStudy (anonymous):

\[(x-1)=4^{3}\]

OpenStudy (anonymous):

\[x=64+1=65\]

OpenStudy (anonymous):

Done x=65

OpenStudy (anonymous):

thanks(: youve been great help

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

do you know how to do (3rdroot of -27)^2